OpenStudy (anonymous):

What are the solutions of 8x2 = 6 + 22x? Check all of the boxes that apply. () x = -3 () x = -1/4 () x = 3 () x = 6

OpenStudy (anonymous):

is that x or multiplication?

OpenStudy (anonymous):

an x

OpenStudy (anonymous):

x= 3

OpenStudy (anonymous):

what other one? theres supposed to be 2 correct choices

OpenStudy (anonymous):

Solve for x over the real numbers: 8 x^2 = 6+22 x Move everything to the left hand side. Subtract 22 x+6 from both sides: -6-22 x+8 x^2 = 0 Factor the left hand side. The left hand side factors into a product with three terms: 2 (-3+x) (1+4 x) = 0 Divide both sides by a constant to simplify the equation. Divide both sides by 2: (-3+x) (1+4 x) = 0 Solve each term in the product separately. Split into two equations: -3+x = 0 or 1+4 x = 0 Look at the first equation: Solve for x. Add 3 to both sides: x = 3 or 1+4 x = 0 Look at the second equation: Isolate terms with x to the left hand side. Subtract 1 from both sides: x = 3 or 4 x = -1 Solve for x. Divide both sides by 4: Answer: | | x = 3 or x = -1/4

OpenStudy (anonymous):

i feel like this website is only being used to cheat haha :)

OpenStudy (anonymous):

Duh ^^

OpenStudy (anonymous):

haha well you pretty much covered it cya

OpenStudy (anonymous):

Where my golden medial!!!

OpenStudy (anonymous):

wow you're like, genius lol thank you hun <3

OpenStudy (anonymous):

Np